OpenStudy (anonymous):

Find the 7th term of the geometric sequence –5, 15, –45, 135, –405,... a. –3,650 b. –3,645 c. 10,935 d. 1,215

OpenStudy (anonymous):

a*r^1=15 I think a*r^6=7th term I think a is -5 use that to find it.

OpenStudy (anonymous):

what about r ? i dont get it. im sorry im too stupid... @timo86m

OpenStudy (anonymous):

U R not the first term is always a so it is -5 in your case the formula is ----------------------- ar^0=1st term , ar^1=2nd term and so on _______________________________ Knowing that 2nd term is 15 and a is -5 we do this -5*r=15 Using algebra we find that r=-3 now we use the formula for the 7th term ar^6 -5*(-3)^6=?

OpenStudy (anonymous):

11390625? @timo86m

OpenStudy (anonymous):

nooo :P

OpenStudy (anonymous):

−3645 @timo86m

OpenStudy (anonymous):

-5*(-3)^6=? Lets leave the -5 outside for now shall we -3*-3*-3*-3*-3*-3=729 now lets drop the -5 in -5*(729) what is that :)

OpenStudy (anonymous):

ooh pretty sure it will be negative 3645

OpenStudy (anonymous):

−3645 !!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

oh sorry here all i see is ??? 3645 :) but anyhoo i am pretty sure -3645

OpenStudy (anonymous):

yes i typed negative but it didnt show thank you so much @timo86m

OpenStudy (anonymous):

that is odd now i think i see them where as before they where ???

OpenStudy (anonymous):

they looked like